OpenStudy (anonymous):

how would I set up this? A road crew repaves 1/28 miles of road a day. How long will it take the crew to repave 4/7 of a mile?

OpenStudy (amistre64):

1/28 = 24 --------- 4/7 = h h = 24(4/7) ------- 1/28

OpenStudy (amistre64):

or to clean it up: h = 24(4)(28) -------- 7

OpenStudy (anonymous):

ty